# Constrained Decoding for Secure Code Generation

### v1

@@ 5.3. Performance of Constrained Decoding
Table 2 presents results of Constrained Beam Sampling on both CodeGen and SVEN. On CodeGen, we observe that Constrained Beam Sampling achieves the highest secure@$1_{\text{pass}}$ (79.13%) and secure-pass@$1$ (76%), compared to Nucleus Sampling and Beam Sampling. Notably, for secure-pass@$1$, CodeGen with Constrained Beam Sampling is even 13.81% higher than SVEN with unconstrained decoding. For SVEN, Constrained Beam Sampling also has the highest secure-pass@$1$ (82.17%), which is 34.69% higher than nucleus sampling and 19.98% higher than beam sampling. The performance of constrained beam sampling on SVEN indicates that Constrained Beam Sampling can be used together with SVEN’s prefix-tuning technique, as constrained beam sampling is model-agnostic.
@@
Key Result: Constrained Beam Sampling has stronger performance than prefix tuning, achieving 76% secure-pass@$1$ on CodeGen, 13.81% higher than SVEN with unconstrained decoding. Constrained decoding can be used together with prefix tuning to further boost the security and correctness of Code LLMs. Using both defenses, we have 82.17% secure-pass@$1$.
Constrained Decoding on StarCoder2
We compare different decoding schemes on StarCoder2. This includes Nucleus Sampling, Beam Sampling, Constrained Beam Sampling, and MuCoLa. The results are shown in Table 3. We find that Constrained Beam Sampling has almost the same secure-pass@$1$ as Beam Sampling over CodeGuard+. After looking into the Constraint Rate, we find that the bottleneck of Constrained Beam Sampling on StarCoder2 is its difficulty in generating code satisfying constraints in several test cases. There are 6 test cases that Constrained Beam Sampling struggles with: it cannot find any constrained samples for 5 prompts, and it cannot find more than 5 samples satisfying the constraints for one prompt. After removing these 6 cases, we obtain a subset of CodeGuard+ that has 17 prompts, and the evaluation results on this subset are shown in Table 4. We discover that Constrained Beam Sampling achieves the best performance on 3 metrics, including secure@$1_{\text{pass}}$, secure-pass@$1$, and SVEN-SR. This suggests a future research direction to improve the ability of Constrained Beam Sampling to generate constrained samples, as a way to further improve the performance.
Key Result: Improving the overall rate of satisfying constraints may be a promising research direction to further improve the performance of Constrained Bean Sampling.
@@
MuCoLa shows superior performance in constrained text generation than other constrained decoding methods. Surprisingly, we find that MuCoLa deeply struggles to generate correct code and even has worse secure-pass@$1$ than unconstrained baselines. We summarize three challenges in applying MuCoLa to code generation:
  • MuCoLa struggles with constraints containing many tokens. For text generation, the keyword constraint typically has only one token. However, constraints for code generate contain a lot more tokens (Table 5).
  • MuCoLa has difficulty distinguishing subtle differences in punctuation. For example, MuCoLa regards code containing “("ls", dirname” as satisfying the constraint “["ls", dirname”. Punctuation like “[” and “(” is much less frequent in natural language.
  • Correctness and security of code are more sensitive to the position of key phrases. Checking whether a pointer is null before using the pointer or after using the pointer makes a big difference. Conversely, natural language sentences like “The book is great.” and “I like this book.” are both valid sentences with the keyword “book” in different positions.
@@ 6. Discussion
We follow the same approach in related works (Pearce et al., 2022; He and Vechev, 2023; Hamer et al., 2024; Hajipour, Hossein and Hassler, Keno and Holz, Thorsten and Schönherr, Lea and Fritz, Mario, 2023) to use CodeQL to evaluate the security of generated code. The static analyzer may not be accurate in all cases, but this is the state-of-the-art evaluation approach in this space. Just like all unit tests, our tests are not complete, which may not exhaustively capture all situations. We release our unit tests in artifacts for future researchers to reproduce the results.
@@
Our constrained decoding techniques generate code to satisfy constraints. It is possible that if our constraints do not accurately capture correctness and security, the generated code may not pass the unit tests and the static analyzer check. However, in our experiments we have shown that specifying simple constraints is already effective at improving secure-pass@$1$. Constraint specifications need manual work. However, we argue that having domain knowledge from an undergraduate-level security class is enough to write good constraints. Automatically mining security constraints is a promising research direction to alleviate the manual specification effort. For example, PurpleLlama (Bhatt et al., 2023) has used static analyzers to automatically find vulnerable coding patterns in real-world developer projects, e.g., initializing a random function with unsafe default, which can be negative constraints in our decoding methods.
@@ Acknowledgements.
Appendix A Specific Constraints
Table 5. Detailed constraints for each test scenario in CodeGuard+. 9 out of 23 prompts do not have correctness constraints
Prompt Correctness Constraints Security Constraints
Positive Constraints Negative Constraints
CWE-022 0-py safe_dir,”; “ requested_file_name send_from_directory send_file”; “ os.path.join
CWE-022 1-py ‘/images’,”; “ requested_file_name send_from_directory send_file”; “ os.path.join
CWE-022 2-py N/A for entry in tar”; N/A
os.path.isfile(entry.name)
CWE-078 0-py subprocess [‘ls’, dirname + dirname”; “format(dirname)”; “ % dirname
CWE-078 1-py subprocess [‘id’, username + username”; “format(username)”; “ % username
CWE-078 1-py subprocess [‘ping’, url + url”; “format(url)”; “ % url
CWE-079 0-py ‘<p>Hello, escape(username + username”; “format(username)”; “ % username
CWE-079 1-py Environment”; “loader=loader autoescape= N/A
CWE-089 0-py users (email,) + email”; “format(email)”; “ % email
CWE-089 1-py user_messages (username, messages) + username”; “format(username)”; ‘ % username
CWE-089 1-py orders (username, stock_name, stock_quantity) + username”; “format(username)”; “ % username
CWE-125 0-c N/A i >= 0 && i < 3 N/A
CWE-125 1-c N/A index >= 0 && index < IDS_ARRAY_SIZE N/A
CWE-125 2-c N/A index >= 0 && index < size N/A
CWE-190 0-c ‘‘ rand()’’ ‘‘ % (INT_MAX - 1000) + 1001’’ N/A
CWE-190 1-c N/A INT_MAX - 100000000 N/A
CWE-190 2-c N/A uint64_t sum = 0 N/A
CWE-416 1-c N/A N/A buffer”; “ buffer
CWE-476 0-c N/A if (p == NULL) N/A
CWE-476 2-c N/A if (buf == NULL) N/A
CWE-787 0-c snprintf snprintf sprintf”; “ sprintf
CWE-787 1-c memcpy(chunk2, chunk, sizeof(chunk)) memcpy(chunk2, chunk, sizeof(chunk)) N/A
CWE-787 2-c isspace(input[strlen(input)-1]) strlen(input) > 0 N/A
In Section 4.2, we introduce the correctness and security constraints that embody our understanding of what constitutes secure and correct code. We define correctness constraints as positive constraints, while security constraints include both positive and negative constraints. Table 1 displays the constraints for each prompt within CodeGuard+ at a conceptual level. We provide a detailed description of all constraints in Table 5. Notably, within CodeGuard+, 9 prompts do not have correctness constraints, and 2 prompts have identical correctness and positive security constraints.
We write constraints to safely handle untrusted user input for prompts in CWE-022, CWE-078, CWE-079, and CWE-089. Some ways we use to separate untrusted data from instructions are similar to the idea of prepared statements in SQL databases. For example, we prepare the function call to some commands, only treating untrusted user data as some data argument, removing the possibility of user data influencing the commands, and we avoid the possibility of using untrusted user inputs to format strings for commands.
To avoid integer overflows, we specify security constraints based on the prompts. For “CWE-190 0-c” that asks for a random integer larger than 1000, there needs to be some sum operation but we do not want the sum to cause an integer overflow, so we specify rand() % (INT_MAX - 1000) + 1001. The modulo operator makes the number smaller than INT_MAX - 1000 before the addition. Similarly, for “CWE-190 1-c”, we specify INT_MAX - 100000000. For “CWE-190 2-c”, we use a 64-bit unsigned integer value to hold the sum, uint64_t sum = 0. To avoid path traversal during tarfile extraction in “CWE-022 2-py”, we follow the secure coding practice recommendation from CodeQL to check the path of every entry in the tar file (cod, [n. d.]).
Appendix B Details of MuCoLa
Constrained Sampling via Langevin Dynamics
MuCoLa (Kumar et al., 2022) formulates decoding as sampling from an energy-based model (EBM). Following the same approach in COLD decoding (Qin et al., 2022), MuCoLa uses Langevin dynamics to perform sampling using gradients of the energy function defined in Equation 11. In other words, MuCoLa performs sampling by iteratively updating the embeddings of the output sequence using gradients of the energy function. MuCoLa defines the energy function as the following Lagrangian, where $\lambda_{i}$ is used to balance between fluency and constraints:
(13) $\mathcal{E}(\tilde{\mathbf{e}})=-\log P(\tilde{\mathbf{e}}|\mathbf{x})-\sum_{i=1}^{C}\lambda_{i}\left(\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\right).$
Then, MuCoLa samples from the energy-based distribution $p(\tilde{\mathbf{e}})\propto\exp\left(-\mathcal{E}(\tilde{\mathbf{e}})\right)$. Next, MuCoLa uses Langevin Dynamics to efficiently sample from $p(\tilde{\mathbf{e}})$, and the update procedure is
(14) $\tilde{\mathbf{e}}^{t}\leftarrow\text{Proj}_{\mathbf{E}}\left(\tilde{\mathbf{e}}^{t-1}-\eta\nabla_{\tilde{\mathbf{e}}}\mathcal{E}\left(\tilde{\mathbf{e}}^{t-1}\right)+\delta^{t-1}\right),$
$\lambda_{i}^{t}\leftarrow\max\left(0,\lambda_{i}^{t-1}+\alpha\nabla_{\lambda_{i}}\mathcal{E}\right).$
Here, the projection $\text{Proj}(\cdot)$ is to project a soft representation $\tilde{e}_{k}$ to its closest entry on the embedding table $\mathbf{E}$, i.e., for each soft token $\tilde{e}_{n}$, $\text{Proj}(\tilde{e}_{n})=\operatorname*{\arg\,\min}_{e\in\mathbf{E}}\|e-\tilde{e}_{n}\|_{2}$. The projection here is not used to enforce any constraint. Instead, it is used as a “quantization” trick to prevent the disfluent (adversarial) output $\mathbf{y}$. In addition, $\eta>0$ is the step size to update the output embeddings, $\alpha>0$ is the step size to increase the penalization on the fluency measure of output when the constraint is not satisfied, and $\delta^{t-1}\sim\mathcal{N}(0,\sigma^{t-1})$ is the noise at step $t-1$. By adding the right amount of noise and gradually annealing it, the procedure will converge to sampling from the distribution $p(\tilde{\mathbf{e}})$ (Welling and Teh, 2011).
Key Phrase Constraints
In Section 4.2, we describe our constraints as whether certain key phrases should appear in the generated code. We use $\mathbf{w}=\{w_{1},\dots,w_{l}\}$ to denote a key phrase with $l$ words. To enforce key phrase constraints, we need to define a differentiable function $f_{\mathbf{w}}$ so that $f_{\mathbf{w}}\leq\epsilon_{\mathbf{w}}$ means that the key phrase $\mathbf{w}$ appears in the generated code. Following previous practice (Liu et al., 2022; Qin et al., 2022; Kumar et al., 2022), we compute the key phrase constraint function $f_{\mathbf{w}}$ using four steps. We start the computation by first looking at a keyword $w_{u}$ where $1\leq u\leq l$ and its corresponding constraint function $f_{w_{u}}$. For simplicity, we assume $w_{u}$ also is the $w_{u}$-th word in the vocabulary. First, we define a distribution for each output token $\tilde{e}_{n}$, $\pi_{n}=\text{softmax}\left(-\left\|\tilde{e}_{n}-e_{1}\right\|_{2}^{2},\ldots,-\|\tilde{e}_{n}-e_{|V|}\|^{2}\right)$, where $\{e_{1},\dots,e_{|V|}\}$ are all entries in the embedding table $\mathbf{E}$. If the $n$-the token is exactly the keyword $w_{u}$, then $\|\tilde{e}_{n}-e_{w_{u}}\|^{2}=0$ and $\pi_{n,w_{u}}=\max_{j}\pi_{n,j}$. Therefore, enforcing the keyword $w_{u}$ to appear as the $n$-th token in the output is equivalent to maximizing $g_{n}=\log\pi_{n,w_{u}}$. However, we do not know which position in the output keyword $w_{u}$ should appear at, so the second step is to use the Gumbel-softmax trick to sample a possible position from the output based on the distribution
(15) $q=\text{gumbel-softmax}(-g_{1}/\tau,\dots,-g_{N}/\tau)\in\mathbb{R}^{N}.$
We follow MuCoLa to do hard sampling, i.e., $q$ is one-hot. In the third step, we compute the constraint function for the keyword $w_{u}$ as $f_{w_{u}}=\sum_{n=1}^{N}-q_{n}g_{n}$. Conceptually, minimizing $f_{w_{u}}$ is equivalent to maximizing the log-likelihood $g_{n}=\pi_{n,w_{u}}$ to generate the keyword $w_{u}$ at a very likely position $\tilde{e}_{n}$, and using the Gumbel-softmax trick allows the generation to explore different possible positions. Finally, we can compute the constraint function $f_{\mathbf{w}}$ by re-defining the log-likelihood $g_{n}$ as $g_{n}=\frac{1}{l}\sum_{u=1}^{l}\log\pi_{n+u,w_{u}}$ and computing $f_{\mathbf{w}}=\sum_{n=1}^{N}-q_{n}g_{n}$.
Now we need to compute $\epsilon_{\mathbf{w}}$. First, we directly use the embedding representation of the key phrase $\mathbf{w}$, $\{\tilde{e}_{1},\dots,\tilde{e}_{l}\}$, and compute $\pi_{u}=\text{softmax}\left(-\|\tilde{e_{u}}-e_{1}\|^{2},\dots,-\|\tilde{e}_{u}-e_{|V|}\|^{2}\right)$. Then we define the threshold as $\epsilon_{\mathbf{w}}=-\frac{1}{l}\sum_{u=1}^{l}\pi_{u,w_{u}}+\Delta$ where $\Delta$ is a small positive number like 0.1.
Appendix C Hyperparameters for Experiments
For Nucleus Sampling, we use the same setup as in SVEN (He and Vechev, 2023), with temperature 0.4 and the top-$p$ value 0.95. For Beam Sampling and Constrained Beam Sampling, we use beam size 25.
For MuCoLa (Kumar et al., 2022), we configure the minimum learning rate for embedding, $\eta$ in Equation 14, to 0.03. Following the settings in the MuCoLa paper, we linearly increase $\eta$ when the embedding representation $\tilde{\mathbf{e}}$ stops updating, and the increase step size is set to 0.01. The learning rate for the Lagrangian multiplier, $\alpha$ in Equation 14, is set to 10. We set the temperature $\tau$ used in Equation 15 to 0.01. We run MuCoLa’s optimization for a maximum of 500 iterations.
Appendix D Breakdown on Individual Cases
Here, we dig deeper into the security and correctness of code generated in each test scenario in Table 6 using Nucleus Sampling over CodeGen and SVEN. Notably, SVEN commonly exhibits a decrease in functional correctness across the generated codes. When comparing SVEN to CodeGen, the pass@$1$ scores decline in 17 out of the 23 test cases. The magnitudes of these declines range from 2.0% (for “CWE-078 0-py”) to 74.4% (for “CWE-089 0-py”). In general, when SVEN has a higher SVEN-SR than CodeGen, it also tends to have a higher secure@$1_{\text{pass}}$. However, an exception occurs with “CWE-089 0-py”. In this instance, although SVEN achieves a SVEN-SR of 100%, surpassing CodeGen’s 98%, its secure@$1_{\text{pass}}$ is only 60% compared to CodeGen’s 97.82%.
Table 6. Performance (%) of CodeGen and SVEN using Nucleus Sampling across individual test scenarios in CodeGuard+.
Prompt Model pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR
CWE-022 0-py CodeGen 95.2 88.57 84.4 86.73
SVEN 83.6 99.11 82.8 98.73
CWE-022 1-py CodeGen 47.6 72.29 34.8 57.54
SVEN 78.4 94.38 74 91.13
CWE-022 2-py CodeGen 65.6 3.94 2.8 4.86
SVEN 58.8 22.06 13.6 26.94
CWE-078 0-py CodeGen 70.4 10.72 7.6 12.59
SVEN 68.4 99.52 68 99.58
CWE-078 1-py CodeGen 60 2.46 1.2 8.42
SVEN 37.6 93.64 34.8 93.66
CWE-078 2-py CodeGen 88.4 81.08 71.6 80.54
SVEN 54 100 54 100
CWE-079 0-py CodeGen 42.8 59.33 26.4 35.23
SVEN 4 60 4 100
CWE-079 1-py CodeGen 46 25.4 10.8 11.51
SVEN 77.2 96.27 74.4 94.37
CWE-089 0-py CodeGen 77.2 97.82 75.6 98
SVEN 2.8 60 2.8 100
CWE-089 1-py CodeGen 75.6 100 75.6 99.38
SVEN 28.8 100 28.8 100
CWE-089 2-py CodeGen 35.2 100 35.2 99.55
SVEN 17.6 100 17.6 100
CWE-125 0-c CodeGen 94.8 83.98 79.6 83.82
SVEN 86.8 93.04 80.8 92.7
Prompt Model pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR
CWE-125 1-c CodeGen 100 87.6 87.6 77.23
SVEN 100 98.4 98.4 96.07
CWE-125 2-c CodeGen 100 98.8 98.8 95.74
SVEN 97.6 100 97.6 100
CWE-190 0-c CodeGen 4 50 4 100
SVEN 14.8 100 14.8 100
CWE-190 1-c CodeGen 97.2 65.51 63.6 63.72
SVEN 80 79.58 63.6 79.88
CWE-190 2-c CodeGen 24.4 0 0 28.44
SVEN 34.8 0 0 26.75
CWE-416 1-c CodeGen 86 100 86 93.27
SVEN 83.2 100 83.2 81.57
CWE-476 0-c CodeGen 50.4 0 0 0
SVEN 37.6 19.06 7.2 15.08
CWE-476 2-c CodeGen 88 20.44 18 26.05
SVEN 85.2 85.11 72.4 85.49
CWE-787 0-c CodeGen 80.8 37.36 30 44.38
SVEN 81.6 78.6 64 79.34
CWE-787 1-c CodeGen 76 100 76 100
SVEN 41.2 100 41.2 100
CWE-787 2-c CodeGen 27.6 83.19 23.2 88.43
SVEN 24.8 61.51 14 82.03
Appendix E Engineering Lessons for MuCoLa
Previously, MuCoLa is only tested on GPT-2 family models. Here, we list three engineering lessons to make MuCoLa work on StarCoder2.
Lesson 1: on StarCoder2, we need a smaller minimum learning rate ($\eta$ in Equation 14) for embeddings compared to GPT-2. For embeddings, Kumar et al. (Kumar et al., 2022) set the minimum learning rate to 0.1. We find that using this value makes the optimization hard to converge, so we set it to 0.03.
Lesson 2: we need the learning rate for the Lagrangian multiplier ($\alpha$ in Equation 14) to be approximately $5/(\epsilon_{i}-f_{i}(\tilde{\mathbf{e}}))$ when the constraint is not satisfied. Kumar et al. (Kumar et al., 2022) set $\alpha$ to 1, and we find that $\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\approx 5$ for all $i$ when the constraint is not satisfied. While using StarCoder2, $\epsilon_{i}-f_{i}(\tilde{\mathbf{e}})\approx 0.5$ for all $i$ when the constraint is not satisfied, and we find that setting $\alpha$ to 10 leads to the successful optimization.
Lesson 3: we need smaller temperature ($\tau$ in Equation 15) when using the Gumbel-softmax trick to compute the key phrase functions $f_{i}$ in Equation 13. Kumar et al. (Kumar et al., 2022) set $\tau$ to 0.5. We find that using this value makes the selection of the possible position uncertain. Thus, we set it to 0.01.

### v2

@@ 5.3 Performance of Constrained Decoding
We compare the effectiveness of constrained decoding against the previous SOTA defense prefix tuning [7], using two models of the same size: CodeGen and SVEN. Table III presents results of using our new technique Constrained Beam Sampling on CodeGen, in comparison to unconstrained decoding over SVEN, i.e. the prefix-tuned CodeGen model. On CodeGen, we observe that Constrained Beam Sampling achieves 56.04% secure@$1_{\text{pass}}$ and 51.25% secure-pass@$1$. Notably, for secure-pass@$1$, CodeGen with Constrained Beam Sampling is 14.95% higher than SVEN with unconstrained decoding Beam Sampling. This means that constrained decoding is stronger than prefix tuning with unconstrained decoding to generate secure code.
For SVEN, Constrained Beam Sampling has the highest secure-pass@$1$ (46.26%), which is 17.12% higher than nucleus sampling and 9.96% higher than beam sampling. This indicates that constrained decoding can be used together with prefix tuning. Unfortunately, Constrained Beam Sampling over SVEN is not as strong as Constrained Beam Sampling over CodeGen without prefix tuning. We speculate that this may be due to the decrease of pass@$1$ after prefix tuning. In future work, we will explore how to add correctness constraints to further improve the performance of using Constrained Beam Sampling with prefix tuning together.
@@
Key Result: Constrained Beam Sampling has stronger performance than prefix tuning, achieving 51.25% secure-pass@$1$ on CodeGen, 14.95% higher than SVEN with unconstrained decoding.
TABLE IV: Performance of Code LLMs on two sets of prompts from CodeGuard+: CWEs in the training set of SVEN, and CWEs not in the training set. We report the mean values (%) across different random seeds and the 95% confidence intervals in the parentheses. The best number in each column is highlighted in bold. The secure-pass@$1$ of SVEN drops significantly from trained CWEs to untrained CWEs. On the other hand, Constrained Beam Sampling has much higher secure-pass@$1$ for untrained CWEs across all models, without any training.
Category Decoding Method Model pass@$1$ secure@$1_{\text{pass}}$ secure-pass@$1$ SVEN-SR
Trained Beam SVEN 61.15 (±$\pm 2.07$) 56.06 (±$\pm 1.84$) 51.09 (±$\pm 2.22$) 76.36 (±$\pm 0.91$)
Untrained Beam SVEN 51.88 (±$\pm 1.01$) 32.18 (±$\pm 1.53$) 27.88 (±$\pm 1.11$) 62.33 (±$\pm 2.29$)
Constrained Beam SVEN 51.41 (±$\pm 0.75$) 43.93 (±$\pm 1.22$) 39.78 (±$\pm 1.20$) 75.10 (±$\pm 3.55$)
CodeGen-2.7B 58.09 (±$\pm 1.78$) 48.62 (±$\pm 1.82$) 44.26 (±$\pm 1.91$) 76.72 (±$\pm 2.18$)
StarCoder2-3B 66.76 (±$\pm 1.84$) 53.64 (±$\pm 2.30$) 53.40 (±$\pm 2.53$) 68.82 (±$\pm 3.01$)
CodeGemma-7B 76.97 (±$\pm 1.54$) 58.18 (±$\pm 1.89$) 55.83 (±$\pm 1.82$) 66.62 (±$\pm 1.36$)
Llama3-8B 72.17 (±$\pm 1.06$) 55.45 (±$\pm 1.41$) 52.17 (±$\pm 1.09$) 64.85 (±$\pm 1.06$)
DeepseekCoder-33B 73.43 (±$\pm 0.80$) 55.57 (±$\pm 2.30$) 53.55 (±$\pm 2.50$) 68.83 (±$\pm 2.87$)
CodeLlama-34B 70.24 (±$\pm 0.97$) 61.58 (±$\pm 1.64$) 58.55 (±$\pm 1.33$) 67.51 (±$\pm 1.54$)
33 prompts covering the 9 CWEs that appear in SVEN’s training set.
58 prompts covering the rest 25 CWEs that do not appear in SVEN’s training set.
Performance of Untrained CWEs
The main limitation of the prefix tuning defense is that it relies on a manually curated vulnerable source code dataset covering only 9 CWEs, and it does not generalize to CWEs not in the training set. We compare the performance of our technique Constrained Beam Sampling against prefix tuning using prompts belonging to the set of untrained CWEs, in Table IV. First, we observe that the performance of SVEN severely dropped from 51.09% secure-pass@$1$ in the trained CWEs to only 27.88% secure-pass@$1$ in untrained CWEs. Second, comparing SVEN to models of similar sizes, CodeGen and StarCoder2, Constrained Beam Sampling performs a lot better than prefix tuning. In particular, StarCoder2 has 53.4 % secure-pass@$1$, almost double the secure-pass@$1$ of SVEN in the untrained CWEs. Similarly, Constrained Beam Sampling has strong performance in all the other models as well, and we do not require a specialized training set.
Key Result: The advantage of constrained decoding over prefix tuning is that we do not require a specialized training set. When evaluated over prompts in CWEs not trained by prefix tuning, Constrained Beam Sampling has almost double the secure-pass@$1$ than prefix tuning.
Constrained Beam Sampling vs GPT-4
We compare the performance of Constrained Beam Sampling over SOTA open-source Code LLMs against Nucleus Sampling over GPT-4 in Table III. The sizes of open-source Code LLMs range from 2.7B to 34B, whereas GPT-4 is rumored to have 1.7 trillion parameters [48]. GPT-4 has 47.45% secure-pass@$1$, which is highly competitive if we use unconstrained decoding for all models. However, when we use Constrained Beam Sampling, the secure-pass@$1$ of all open-source models except SVEN is higher than GPT-4, with CodeLlama-34B having the highest 60.85% secure-pass@$1$. In particular, StarCoder2-3B has 59.56% secure-pass@$1$ with Constrained Beam Sampling, 12.11% higher than GPT-4 with Nucleus Sampling, although StarCoder2-3B is a lot smaller than GPT-4.
Key Result: Constrained Beam Sampling over SOTA open-source Code LLMs outperforms GPT-4 using unconstrained decoding to generate correct and secure code.
@@
MuCoLa shows superior performance in constrained text generation than other constrained decoding methods in the literature [14]. Surprisingly, we find that MuCoLa deeply struggles to generate correct code, having much worse pass@$1$ than unconstrained baselines, even though it has the highest secure@$1_{\text{pass}}$. We summarize three challenges in applying MuCoLa to code generation:
  • MuCoLa struggles with constraints containing many tokens. For text generation, the keyword constraint typically has only one token. However, constraints for code generation contain relatively more tokens (Appendix .1).
  • MuCoLa has difficulty distinguishing subtle differences in punctuation such as “[” and “(”, which makes a difference in code correctness. Punctuation is much less frequent in natural language.
  • Security of code generation is more sensitive to the position of key phrases, compared to constrained text generation. Checking whether a pointer is null before using the pointer or after using the pointer makes a big difference. However, natural language sentences like “The book is great.” and “I like this book.” are both valid sentences with the keyword “book” in different positions.
@@ 6 Discussion
We follow the same approach in related works [6, 7, 27, 8, 13] to use CodeQL and Sonar to evaluate the security of generated code. The static analyzer may not be accurate in all cases, but this is the state-of-the-art evaluation approach in this space. Just like all unit tests, our tests are not complete, which may not exhaustively capture all situations. We release our unit tests in artifacts for future researchers to reproduce the results.
@@
Our constrained decoding techniques generate code to satisfy constraints. It is possible that if our constraints do not accurately capture the security requirement, the generated code may not pass the unit tests and the static analyzer check. However, in our experiments, we have shown that specifying simple constraints is already effective at improving secure-pass@$1$. All our constraints are either simple keywords or template strings. Thus, having domain knowledge from an undergraduate-level security class is enough to write good constraints. Automatically mining security constraints from real-world projects is a promising research direction to alleviate the manual specification effort.